# ChartMimic: Evaluating LMM’s Cross-Modal Reasoning Capability via Chart-to-Code Generation

### v1

@@ Appendix B Chart Taxonomy
Appendix C Details of Evaluation Metrics
@@
In this section, we present the details of evaluation metrics, including GPT-4V score for high-level metrics and Text Score, Layout Score, Type Score and Color Score for low-level metrics. For more information, please refer to our source code.
C.1 GPT-4V Score
@@
We employ GPT-4V [2] to assess the extend to which the generated figure corresponds to the ground-truth figure. The specific content of the prompt is presented in Fig. 27. Specifically, we input both the generated and the ground-truth figures into the GPT-4V simultaneously. Then, GPT-4V is instructed to evaluate the similarity between the two figures, taking into account six dimensions: text, layout, type, data, style, and clarity. Subsequently, GPT-4V outputs a score ranging from 0 to 100 to represent the degree of similarity between the figures.
@@
Prompt for GPT-4V Score You are an excellent judge at evaluating visualization chart plots. The first image (reference image) is created using ground truth matplotlib code, and the second image (AI-generated image) is created using matplotlib code generated by an AI assistant. Your task is to score how well the AI-generated plot matches the ground truth plot. ### Scoring Methodology: The AI-generated image’s score is based on the following criteria, totaling a score out of 100 points: 1. Chart Types (20 points): Does the AI-generated image include all chart types present in the reference image (e.g., line charts, bar charts, etc.)? 2. Layout (10 points): Does the arrangement of subplots in the AI-generated image match the reference image (e.g., number of rows and columns)? 3. Text Content (20 points): Does the AI-generated image include all text from the reference image (e.g., titles, annotations, axis labels), excluding axis tick labels? 4. Data (20 points): How accurately do the data trends in the AI-generated image resemble those in the original image and is the number of data groups the same as in the reference image? 5. Style (20 points): Does the AI-generated image match the original in terms of colors (line colors, fill colors, etc.), marker types (point shapes, line styles, etc.), legends, grids, and other stylistic details? 6. Clarity (10 points): Is the AI-generated image clear and free of overlapping elements? ### Evaluation: Compare the two images head to head and provide a detailed assessment. Use the following format for your response: Comments: - Chart Types: \${your comment and subscore} - Layout: \${your comment and subscore} - Text Content: \${your comment and subscore} - Data: \${your comment and subscore} - Style: \${your comment and subscore} - Clarity: \${your comment and subscore} Score: \${your final score out of 100} Please use the above format to ensure the evaluation is clear and comprehensive.
Figure 27: Prompt for GPT-4V Score.
C.2 Text Score
@@
In order to accurately capture the textual content presented in the rendered figures, our code tracer monitors the function used to add text to the rendered PDF, logging each textual element. An exemplary Python code is provided in Listing LABEL:lst:text, where we record the text elements by adding a log wrapper to the “draw_text()” function in the matplotlib package. Subsequently, we employ this approach to extract the text information from both the generated code and the ground-truth code. The F1 score, computed between these two groups of texts, is denoted as Text Score.
C.3 Layout Score
@@
The layout refers to the organization of subplots within a chart figure. In each figure implemented using matplotlib, multiple axes objects are present, each containing layout information that specifies its position within the figure. To analyze the layout, we iterate through each axis object in the figure and obtain their respective layout information. An exemplary Python code to accomplish this process is provided in Listing LABEL:lst:layout, where the position information of each axis is recorded. Subsequently, we gather the layout information from both the ground-truth code and the generated code, and compute the F1 score between the them. This metric is referred as the Layout Score.
C.4 Type Score
@@
The matplotlib package provides a variety of functions for easily generating diverse chart types, such as using “axes.bar()” to create bar charts. By monitoring the invocations of these functions, we can identify the types of charts being utilized. A successful invocation of a plot functions indicates the incorporation of a specific chart type in the final visualization. Listing LABEL:lst:charttype presents an exemplary Python code that demonstrates this approach, where we introduce a logger to the “bar()” function. Subsequently, we gather the chart types from both the generated and ground-truth code, and compute the F1 score between them. We refer to this metric as Type Score.
C.5 Color Score
@@
In the matplotlib package, each plot function returns a chart type instance at the end of the function invocation. These instances contain various attributes, including those related to color properties, such as facecolor, edgecolor and colormap. To assess the color attributes, we employ the code tracer that captures the color information of each chart type instance. An example Python code demonstrating this approach is provided in Listing LABEL:lst:color. Once we obtain the color attributes from both the ground-truth code and the generated code, we compare them using the F1 score metric.
It is noteworthy that through preliminary experiments, we find that using exact color matching resulted in very low color similarity, as even slight variations would result in a similarity score of zero. To address this issue, we employ the CIEDE2000 color difference formula[51], which converts the matching value between two colors from a discrete [0,1] scale to a continuous range between 0 and 1. Finally, we calculate the maximum color similarity between the sets of colors in the ground-truth code and the generated code to compute F1 score. We refer to the F1 score as Color Score.
Appendix D Generation Configurations and Prompting Methods
D.1 Generation Configurations
Following previous setups [52, 53], for open-weight models, we set the temperature $\tau=0.1$ to achieve optimal results, while for proprietary models, we set the temperature $\tau=0$ for greedy decoding. For all models, we set the maximum generation length to 4096. Additionally, we use BF16 for model inference for open-weight models. All models are inferred on a single A100 80G GPU.
D.2 Prompts
We provide prompts for Direct, HintEnhanced, SelfRevision and Scaffold Prompting in Figs. 28, 29, 30, 31 and 32. for Direct Prompting, we meticulously design a separate prompt for open-source models to achieve optimal results, as shown inFig. 28. while the prompt for proprietary Models is shown in Fig. 29.
Prompt for Direct Prompting (Open-Weight Models) You are an expert Python developer who specializes in writing matplotlib code based on a given picture. I found a very nice picture in a STEM paper, but there is no corresponding source code available. I need your help to generate the Python code that can reproduce the picture based on the picture I provide. Please note that it is necessary to use figsize=({width} (width), {height} (height)) to set the image size to match the original size. Additionally, I will not provide you with the actual data in the image, so you have to extract the actual data by yourself and based on the extracted data to reproduce the image. Ensure that the code you provide can be executed directly without requiring me to add additional variables. Now, please give me the matplotlib code that reproduces the picture below.
@@ D.2 Prompts
Prompt for Direct Prompting (Proprietary Models) You are an expert Python developer who specializes in writing matplotlib code based on a given picture. I found a very nice picture in a STEM paper, but there is no corresponding source code available. I need your help to generate the Python code that can reproduce the picture based on the picture I provide. Note that it is necessary to use figsize=({width} (width), {height} (height)) to set the image size to match the original size. Now, please give me the matplotlib code that reproduces the picture below.
@@
Prompt for HintEnhanced Prompting You are an expert Python developer who specializes in writing matplotlib code based on a given picture. I found a very nice picture in a STEM paper, but there is no corresponding source code available. I need your help to generate the Python code that can reproduce the picture based on the picture I provide. To ensure accuracy and detail in your recreation, begin with a comprehensive analysis of the figure to develop an elaborate caption. This caption should cover, but not be limited to, the following aspects: 1. Layout Analysis: e.g., identify the picture’s composition, noting the presence and arrangement of any subplots. 2. Chart Type Identification: e.g., determine how many charts within a subplot. Are they independent, or do they share a common axis? 3. Data Analysis: e.g., summarize the data trend or pattern. 4. Additional Features: e.g.,identify any supplementary elements such as legends, colormaps, tick labels, or text annotations that contribute to the figure’s clarity or aesthetic appeal. Now, given the picture below, please first output your comprehensive caption and then use the caption to assist yourself to generate matplotlib code that reproduces the picture. Note that it is necessary to use figsize=({width} (width), {height} (height)) to set the image size to match the original size.
@@
Prompt for SelfRevision Prompting You are an expert Python developer who specializes in writing matplotlib code based on a given picture. I have a code for implementing the reference picture as follows: Now, I have the Python matplotlib code for implementing the reference picture as follows: ```python {python_function} ``` The rendered picture of the code is: Now, please compare whether the renderer picture is the same as the reference picture. The difference may cover, but not be limited to, the following aspects: 1. Chart Types: Does the AI-generated image include all chart types present in the reference image (e.g., line charts, bar charts, etc.)? 2. Layout: Does the arrangement of subplots in the AI-generated image match the reference image (e.g., number of rows and columns)? 3. Text Content: Does the AI-generated image include all text from the reference image (e.g., titles, annotations, axis labels), excluding axis tick labels? 4. Data: How accurately do the data trends in the AI-generated image resemble those in the original image and is the number of data groups the same as in the reference image? 5. Style: Does the AI-generated image match the original in terms of colors (line colors, fill colors, etc.), marker types (point shapes, line styles, etc.), legends, grids, and other stylistic details? 6. Clarity: Is the AI-generated image clear and free of overlapping elements? - If the generated picture matches the reference, please output the original implementation code. - If there are discrepancies, first list the specific differences between the two pictures. Then, modify the existing code to address these differences, ensuring the revised code is capable of reproducing the reference picture. Finally, output the revised code. Note that it is necessary to use figsize=({width} (width), {height} (height)) to set the image size to match the original size.
Figure 31: Prompt for SelfRevision Prompting. {text} in blue font represents placeholders, which varies according to different test examples.
Prompt for Scaffold Prompting You are an expert Python developer who specializes in writing matplotlib code based on a given picture. I found a very nice picture in a STEM paper, but there is no corresponding source code available. I need your help to generate the Python code that can reproduce the picture based on the picture I provide. I will provide you with two images. The first image is the original picture. The second image is the picture overlaid with a dot matrix of a shape of {dot_matrix_height} * {dot_matrix_width} to help you with your task, and each dot is labeled with two-dimensional coordinates (x,y). Within each column, the x-coordinate increases from top to bottom, and within each row, the y-coordinate increases from left to right. Please first use this dot matrix as reference anchors to generate the description of the picture (e.g., between dot A and dot B is something) and generate matplotlib code that reproduces the picture. Note that it is necessary to use figsize=({width} (width), {height} (height)) to set the image size to match the original size.
@@
D.3 Error Cases of Different Prompting Methods
We provide error cases of HintEnhanced, SelfRevision and Scaffold Prompting in Figs. 33, 34 and 35. For an analysis of error cases related to Direct Prompting, please refer to Appendix F.
Figure 33: An error case of HintEnhanced Prompting. The text highlighted in blue is GPT-4V’s incorrect understanding regarding provided figures. The remaining part of response is omitted with “……”.
Figure 34: An error case of SelfRevision Prompting. The text highlighted in blue is GPT-4V’s incorrect understanding regarding provided figures. The remaining part of response is omitted with “……”.
Figure 35: An error case of Scaffold Prompting. The text highlighted in blue is GPT-4V’s incorrect understanding regarding provided figures. The remaining part of response is omitted with “……”.
Appendix E Correlation with Human Evaluation

### v2

@@ Appendix D Chart Taxonomy
Appendix E Details of Evaluation Metrics
@@
In this section, we present the details of evaluation metrics, including GPT-4o score for high-level metrics and Text Score, Layout Score, Type Score and Color Score for low-level metrics. For more information, please refer to our source code.
E.1 GPT-4o Score
@@
We employ GPT-4o (OpenAI, 2024) to assess the extend to which the generated figure corresponds to the ground-truth figure. The specific content of the prompt is presented in Fig. 34. Specifically, we input both the generated and the ground-truth figures into the GPT-4o simultaneously. Then, GPT-4o is instructed to evaluate the similarity between the two figures, taking into account six dimensions: text, layout, type, data, style, and clarity. Subsequently, GPT-4o outputs a score ranging from 0 to 100 to represent the degree of similarity between the figures.
@@
Prompt for GPT-4o Score You are an excellent judge at evaluating visualization chart plots. The first image (reference image) is created using ground truth matplotlib code, and the second image (AI-generated image) is created using matplotlib code generated by an AI assistant. Your task is to score how well the AI-generated plot matches the ground truth plot. ### Scoring Methodology: The AI-generated image’s score is based on the following criteria, totaling a score out of 100 points: 1. Chart Types (20 points): Does the AI-generated image include all chart types present in the reference image (e.g., line charts, bar charts, etc.)? 2. Layout (10 points): Does the arrangement of subplots in the AI-generated image match the reference image (e.g., number of rows and columns)? 3. Text Content (20 points): Does the AI-generated image include all text from the reference image (e.g., titles, annotations, axis labels), excluding axis tick labels? 4. Data (20 points): How accurately do the data trends in the AI-generated image resemble those in the original image and is the number of data groups the same as in the reference image? 5. Style (20 points): Does the AI-generated image match the original in terms of colors (line colors, fill colors, etc.), marker types (point shapes, line styles, etc.), legends, grids, and other stylistic details? 6. Clarity (10 points): Is the AI-generated image clear and free of overlapping elements? ### Evaluation: Compare the two images head to head and provide a detailed assessment. Use the following format for your response: Comments: - Chart Types: \${your comment and subscore} - Layout: \${your comment and subscore} - Text Content: \${your comment and subscore} - Data: \${your comment and subscore} - Style: \${your comment and subscore} - Clarity: \${your comment and subscore} Score: \${your final score out of 100} Please use the above format to ensure the evaluation is clear and comprehensive.
Figure 34: Prompt for GPT-4o Score.
Stability of Evaluation with GPT-4o Score. We conduct the high-level evaluation for GPT-4o on the Direct Mimic task for 5 times to assess the stability of GPT-4o Score. The result indicate a mean GPT-4o Score of $83.4$ with a standard deviation of 0.08, demonstrating the stability of GPT-4o Score.
Cost of Evaluation with GPT-4o Score. A single-round evaluation with GPT-4o Score on the Direct Mimic task approximately costs \$5.25. Utilizing OpenAI’s batch services further reduces this cost to \$2.63 per round.
Longevity of Evaluation with GPT-4o Score. The GPT-4o Score metric is designed with long-term viability in mind. Our approach provides a meta-evaluation framework that can adapt to evolving language models. While currently leveraging GPT-4o, the method is model-agnostic and can be implemented with other advanced LLMs such as Claude or Gemini. This adaptability ensures that as more capable models emerge, the evaluation process can seamlessly transition to utilize these superior systems. Importantly, this approach has demonstrated increasing correlation with human assessments as model capabilities improve, as evidenced by studies like AlpacaEval (Li et al., 2023) and MT-Bench (Zheng et al., 2023).
Figure 35: Example of GPT-4o’s scoring results for a Direct Mimic example (bar_28).
Figure 36: Example of GPT-4o’s scoring results for a Customized Mimic example (bar_28).
Figure 37: Example of GPT-4o’s scoring results for a Direct Mimic example (CB_29).
Figure 38: Example of GPT-4o’s scoring results for a Customized Mimic example (CB_29).
GPT-4o Score Examples. To demonstrate our evaluation framework, we present GPT-4o’s scoring examples for both Direct Mimic and Customized Mimic tasks. Figs. 35, 36, 37 and 38 show the evaluation results for two representative charts of different complexity levels (bar_28 and CB_29). These examples demonstrate how GPT-4o systematically evaluates various aspects of chart reproduction across different chart complexities and task types.
Correlation Coefficient Comparison with CLIP Score. CLIP Score (Radford et al., 2021) is widely used for assessing figure similarity. However, our preliminary experiments indicate that it struggles to differentiate variations in types and other critical elements in charts, resulting in a low correlation coefficient of 0.53 with human evaluation. In contrast, as stated in Sec. 4.3, GPT-4o Score achieves a correlation coefficient of 0.70 with human evaluation. Therefore, we adopt GPT-4o Score as the high-level evaluation metric.
E.2 Text Score
@@
In order to accurately capture the textual content presented in the rendered figures, our code tracer monitors the function used to add text to the rendered PDF, logging each textual element. An exemplary Python code is provided in Listing LABEL:lst:text, where we record the text elements by adding a log wrapper to the “draw_text()” function in the matplotlib package. Subsequently, we employ this approach to extract the text information from both the generated code and the ground-truth code.
Based on these two groups of texts, precision is defined as the ratio of the number of correctly captured ground-truth text to the total number of text in the generated figure. Recall is defined as the ratio of the number of correctly captured ground-truth text to the total number of text in the ground-truth figure. The F1-score, calculated using precision and recall, serves as Text Score.
E.3 Layout Score
@@
The layout refers to the organization of subplots within a chart figure. In each figure implemented using matplotlib, multiple axes objects are present, each containing layout information that specifies its position within the figure. To analyze the layout, we iterate through each axis object in the figure and obtain their respective layout information. An exemplary Python code to accomplish this process is provided in Listing LABEL:lst:layout, where the position information of each axis is recorded. Subsequently, we gather the layout information from both the ground-truth code and the generated code.
Based on these two groups of layout information, precision is defined as the ratio of the number of correctly identified ground-truth layouts to the total number of layouts in the generated figure. Recall is defined as the ratio of the number of correctly identified ground-truth layouts to the total number of layouts in the ground-truth figure. The F1-score, calculated using precision and recall, serves as Layout Score.
E.4 Type Score
@@
The matplotlib package provides a variety of functions for easily generating diverse chart types, such as using “axes.bar()” to create bar charts. By monitoring the invocations of these functions, we can identify the types of charts being utilized. A successful invocation of a plot functions indicates the incorporation of a specific chart type in the final visualization. Listing LABEL:lst:charttype presents an exemplary Python code that demonstrates this approach, where we introduce a logger to the “bar()” function. Subsequently, we gather the chart types from both the generated and ground-truth code.
Based on these two groups of chart types, precision is defined as the ratio of the number of correctly identified ground-truth chart types to the total number of chart types in the generated figure. Recall is defined as the ratio of the number of correctly identified ground-truth chart types to the total number of chart types in the ground-truth figure. The F1-score, calculated using precision and recall, serves as Type Score.
E.5 Color Score
@@
In the matplotlib package, each plot function returns a chart type instance at the end of the function invocation. These instances contain various attributes, including those related to color properties, such as facecolor, edgecolor and colormap. To assess the color attributes, we employ the code tracer that captures the color information of each chart type instance. An example Python code demonstrating this approach is provided in Listing LABEL:lst:color. Subsequently, we gather the color information from both the ground-truth code and the generated code and calculate the similarity between them.
It is noteworthy that through preliminary experiments, we find that using exact color matching resulted in very low color similarity, as even slight variations would result in a similarity score of zero. To address this issue, we employ the CIEDE2000 color difference formula (Luo et al., 2001), which converts the matching value between two colors from a discrete $[0,1]$ scale to a continuous range between 00 and $1$. Finally, we calculate the maximum color similarity between the sets of colors in the ground-truth code and the generated code. Precision is defined as the ratio of maximum color similarity to the total number of color in the generated figure. Recall is defined as the ratio of maximum color similarity to the total number of color in the ground-truth figure. The F1-score, calculated using precision and recall, serves as Color Score.
Appendix F Model Configurations and Prompting Methods
F.1 Generation Configurations
Following previous setups (Wang et al., 2023b; Shi et al., 2024), for open-weight models, we set the temperature $\tau=0.1$ to achieve optimal results, while for proprietary models, we set the temperature $\tau=0$ for greedy decoding. For all models, we set the maximum generation length to 4096. Additionally, we use BF16 for model inference for open-weight models. All models are inferred on A100 80G GPU.
F.2 Prompts
We provide prompts for Direct, HintEnhanced, SelfReflection and Scaffold Prompting in Figs. 39, 40, 41, 43 and 42. for Direct Prompting, we meticulously design a separate prompt for open-source models to achieve optimal results, as shown in Fig. 39. while the prompt for proprietary Models is shown in Fig. 40.
Prompt for Direct Prompting (Open-Weight Models) You are an expert Python developer who specializes in writing matplotlib code based on a given picture. I found a very nice picture in a STEM paper, but there is no corresponding source code available. I need your help to generate the Python code that can reproduce the picture based on the picture I provide. Please note that it is necessary to use figsize=({width}, {height}) to set the image size to match the original size. Additionally, I will not provide you with the actual data in the image, so you have to extract the actual data by yourself and based on the extracted data to reproduce the image. Ensure that the code you provide can be executed directly without requiring me to add additional variables. Now, please give me the matplotlib code that reproduces the picture below.
@@ F.2 Prompts
Prompt for Direct Prompting (Proprietary Models) You are an expert Python developer who specializes in writing matplotlib code based on a given picture. I found a very nice picture in a STEM paper, but there is no corresponding source code available. I need your help to generate the Python code that can reproduce the picture based on the picture I provide. Note that it is necessary to use figsize=({width}, {height}) to set the image size to match the original size. Now, please give me the matplotlib code that reproduces the picture below.
@@
Prompt for HintEnhanced Prompting You are an expert Python developer who specializes in writing matplotlib code based on a given picture. I found a very nice picture in a STEM paper, but there is no corresponding source code available. I need your help to generate the Python code that can reproduce the picture based on the picture I provide. To ensure accuracy and detail in your recreation, begin with a comprehensive analysis of the figure to develop an elaborate caption. This caption should cover, but not be limited to, the following aspects: 1. Layout Analysis: e.g., identify the picture’s composition, noting the presence and arrangement of any subplots. 2. Chart Type Identification: e.g., determine how many charts within a subplot. Are they independent, or do they share a common axis? 3. Data Analysis: e.g., summarize the data trend or pattern. 4. Additional Features: e.g.,identify any supplementary elements such as legends, colormaps, tick labels, or text annotations that contribute to the figure’s clarity or aesthetic appeal. Now, given the picture below, please first output your comprehensive caption and then use the caption to assist yourself to generate matplotlib code that reproduces the picture. Note that it is necessary to use figsize=({width}, {height}) to set the image size to match the original size.
@@
Prompt for Scaffold Prompting You are an expert Python developer who specializes in writing matplotlib code based on a given picture. I found a very nice picture in a STEM paper, but there is no corresponding source code available. I need your help to generate the Python code that can reproduce the picture based on the picture I provide. I will provide you with two images. The first image is the original picture. The second image is the picture overlaid with a dot matrix of a shape of {dot_matrix_height} * {dot_matrix_width} to help you with your task, and each dot is labeled with two-dimensional coordinates (x,y). Within each column, the x-coordinate increases from top to bottom, and within each row, the y-coordinate increases from left to right. Please first use this dot matrix as reference anchors to generate the description of the picture (e.g., between dot A and dot B is something) and generate matplotlib code that reproduces the picture. Note that it is necessary to use figsize=({width}, {height}) to set the image size to match the original size.
@@
Prompt for SelfReflection Prompting You are an expert Python developer who specializes in writing matplotlib code based on a given picture. I have a code for implementing the reference picture as follows: Now, I have the Python matplotlib code for implementing the reference picture as follows: ```python {python_code} ``` The rendered picture of the code is: Now, please compare whether the renderer picture is the same as the reference picture. The difference may cover, but not be limited to, the following aspects: 1. Chart Types: Does the AI-generated image include all chart types present in the reference image (e.g., line charts, bar charts, etc.)? 2. Layout: Does the arrangement of subplots in the AI-generated image match the reference image (e.g., number of rows and columns)? 3. Text Content: Does the AI-generated image include all text from the reference image (e.g., titles, annotations, axis labels), excluding axis tick labels? 4. Data: How accurately do the data trends in the AI-generated image resemble those in the original image and is the number of data groups the same as in the reference image? 5. Style: Does the AI-generated image match the original in terms of colors (line colors, fill colors, etc.), marker types (point shapes, line styles, etc.), legends, grids, and other stylistic details? 6. Clarity: Is the AI-generated image clear and free of overlapping elements? - If the generated picture matches the reference, please output the original implementation code. - If there are discrepancies, first list the specific differences between the two pictures. Then, modify the existing code to address these differences, ensuring the revised code is capable of reproducing the reference picture. Finally, output the revised code. Note that it is necessary to use figsize=({width}, {height}) to set the image size to match the original size.
Figure 43: Prompt for SelfReflection Prompting. {text} in blue font represents placeholders, which varies according to different test examples.
F.3 Cases of Different Prompting Methods
We provide cases of HintEnhanced, SelfReflection and Scaffold Prompting in Figs. 44, 45, 46 and 47. For an analysis of error cases related to Direct Prompting, please refer to Appendix I.
Figure 44: An error case of HintEnhanced Prompting. The text highlighted in red is LMM’s incorrect understanding regarding provided figures. The remaining part of response is omitted with “……”.
Figure 45: An error case of SelfReflection Prompting. The text highlighted in red is LMM’s incorrect understanding regarding provided figures. The remaining part of response is omitted with “……”.
Figure 46: An good case of SelfReflection Prompting. The text highlighted in blue is LMM’s correct reflection regarding provided figures. The remaining part of response is omitted with “……”.
Figure 47: An error case of Scaffold Prompting. The text highlighted in red is LMM’s incorrect understanding regarding provided figures. The remaining part of response is omitted with “……”.
F.4 Details of Models
We list the evaluated models in Tab. 9.
Model Params Language Model Vision Model Model Code/API
GPT-4o - - - gpt-4o-2024-05-13
Claude-3-opus - - - claude-3-opus-20240229
GeminiProVision - - - gemini-pro-vision
InternVL2-Llama3-76B 76.0B Llama-3-70B-Instruct InternViT-6B OpenGVLab/InternVL2-Llama3-76B
LLaVA-Next-Yi-34B 34.8B Yi-34B CLIP ViT-L/14 llava-hf/llava-v1.6-34b-hf
InternVL2-26B 26.0B InternLM2-20B InternViT-6B OpenGVLab/InternVL2-26B
Cogvlm2-llama3-chat-19B 19.2B Llama-3-8B-Instruct EVA2-CLIP-E THUDM/cogvlm2-llama3-chat-19B
Phi-3-Vision-128K-Instruct 4.2B Phi-3 CLIP ViT-L/14 microsoft/Phi-3-vision-128k-instruct
IDEFICS2-8B 7.6B Mistral-7B SigLip-400M HuggingFaceM4/idefics2-8b
LLaVA-Next-Mistral-7B 7.6B Mistral-7B CLIP ViT-L/14 llava-hf/llava-v1.6-mistral-7b-hf
DeepSeek-VL-7B 7.3B DeekSeek-7B SAM-B & SigLIP-L deepseek-ai/deepseek-vl-7b-chat
MiniCPM-Llama3-V2.5 8.4B Llama3-8B-Instruct SigLip-400M openbmb/MiniCPM-Llama3-V-2_5
Qwen2-VL-7B 8.2B Qwen2-7B ViT-600M Qwen/Qwen2-VL-7B-Instruct
InternVL2-8B 8.1B InternLM2.5-7B InternViT-300M OpenGVLab/InternVL2-8B
InternVL2-4B 4.2B Phi-3 InternViT-300M OpenGVLab/InternVL2-4B
Qwen2-VL-2B 2.6B Qwen2-1.5B ViT-600M Qwen/Qwen2-VL-2B-Instruct
InternVL2-2B 2.2B InternLM2-1.8B InternViT-300M OpenGVLab/InternVL2-2B
Table 9: Model code/API of the evaluated models.
Appendix G Correlation with Human Evaluation